OpenStudy (anonymous):

I truly need help with systems of Equations! 1. -2x+3y=8 3x-y=-5

OpenStudy (mrnood):

Can you re-arrange 3x-y=5 into y= ????

OpenStudy (anonymous):

Multiply the second expression by 3, so it becomes 9x-3y=-15 Add first expression to second -2x+9x+3y-3y=8-15 You should get 7x=-7 after simplifying the top expression The system of equations now should look like this 7x=-7 3x-y=-5 Find the value of x from the first expression x=-7/7=-1 So x=-1 3x-y=-5 Substitute the value of x in the second expression to get -3-y=-5, rearrange the expression against y y=5-3=2 You final answers x=-1; y=2.

OpenStudy (imstuck):

Probably the best way to do this would be to use the addition method. That is where you have to multiply both equations by a specific number to make either the x terms or the y terms be the same number with opposite signs. Like 2x and -2x cancel each other out, right? Just an example. So here we have a 3y and a -y. In order to make the second equation's y term be a -3y to cancel with the positive 3y, we need to multiply the -y by a 3. 3(-y) = -3y. But if you multiply the -y by a 3, you have to multiply the whole equation by a 3. So here's what you would have if you do that.\[-2x+3y=8\] and\[3(3x-y=-5)->9x-3y=-15\]